# Benchmarking Open-ended Audio Dialogue Understanding for Large Audio-Language Models

### v1

@@ Abstract
Large Audio-Language Models (LALMs) have unclocked audio dialogue capabilities, where audio dialogues are a direct exchange of spoken language between LALMs and humans. Recent advances, such as GPT-4o, have enabled LALMs in back-and-forth audio dialogues with humans. This progression not only underscores the potential of LALMs but also broadens their applicability across a wide range of practical scenarios supported by audio dialogues. However, given these advancements, a comprehensive benchmark to evaluate the performance of LALMs in the open-ended audio dialogue understanding remains absent currently. To address this gap, we propose an Audio Dialogue Understanding Benchmark (ADU-Bench), which consists of 4 benchmark datasets. They assess the open-ended audio dialogue ability for LALMs in 3 general scenarios, 12 skills, 9 multilingual languages, and 4 categories of ambiguity handling. Notably, we firstly propose the evaluation of ambiguity handling in audio dialogues that expresses different intentions beyond the same literal meaning of sentences, e.g., “Really!?” with different intonations. In summary, ADU-Bench includes over 20,000 open-ended audio dialogues for the assessment of LALMs. Through extensive experiments conducted on 13 LALMs, our analysis reveals that there is still considerable room for improvement in the audio dialogue understanding abilities of existing LALMs. In particular, they struggle with mathematical symbols and formulas, understanding human behavior such as roleplay, comprehending multiple languages, and handling audio dialogue ambiguities from different phonetic elements, such as intonations, pause positions, and homophones.
@@ 1 Introduction
Large Audio-Language Models (LALMs) (Zhang et al., 2023; Wang et al., 2023; Chu et al., 2023; Tang et al., 2024; Wu et al., 2023; Kong et al., 2024; Lin et al., 2024; Xie and Wu, 2024; Fu et al., 2024) have received attention for their abilities to handle various audio-related tasks. In particular, LALMs recently unlock unprecedented capabilities for interactive audio dialogues with humans. These dialogues are defined as a direct exchange of spoken language between LALMs and humans, which fosters a more engaging and dynamic mode of communication. Recent advances, such as GPT-4o (OpenAI, 2024), have enabled LALMs to engage in back-and-forth dialogues with humans and can observe various audio characteristics, which broadens their applicability across diverse real-world situations that rely on interactive audio dialogues.
@@
However, given these advancements, there is currently no comprehensive benchmark to evaluate LALMs’ performance in handling open-ended audio dialogue understanding. Previous benchmarks on LALMs predominantly focus on their performance in multiple fundamental tasks, such as speech-to-text translation (Wang et al., 2020a, b; Jia et al., 2022), emotion recognition (Cao et al., 2014; Livingstone and Russo, 2018; Yang et al., 2024), and audio question answering with textual prompts (Yang et al., 2024), etc. While these tasks are essential, they do not adequately capture the diversity inherent in real-world audio dialogues. The absence of a comprehensive benchmark for evaluating LALMs in open-ended audio dialogues has led to suboptimal comparisons between different LALMs. This gap in evaluation benchmarks hinders the development of existing LALMs.
Furthermore, open-ended audio dialogues, where users can directly engage with LALMs through audio, constitute a significant portion of real-world interactions. These dialogues can encompass many subjects, such as helpful and daily questions, domain-specific skills, and multiple different languages. Additionally, the variations in intonations or pause positions can allow speakers to express different intentions beyond the same literal meaning of sentences, adding further complexity to the dialogues. Therefore, the ability to handle open-ended audio dialogues effectively is crucial for LALMs to be truly useful in real-world applications. In light of these considerations, there is a pressing need to establish a benchmark that can effectively assess the performance of LALMs in handling these challenges of audio dialogues.
In this work, we propose an Audio Dialogue Understanding Benchmark (ADU-Bench), a benchmark to evaluate the open-ended audio dialogue understanding for LALMs, which comprises 4 benchmark datasets as follows. (1) The ADU-General dataset assesses the general dialogue understanding of LALMs, including 3 scenarios, i.e., helpful questions to query search engines, daily questions happening among human dialogues, and daily statements without rich contexts. (2) The ADU-Skill dataset evaluates the skill-based dialogue ability, encompassing 12 different skills such as mathematics, physics, coding, etc. (3) The ADU-Multilingual dataset tests the multilingual dialogue understanding, covering 9 languages, including English, French, and Chinese, etc. (4) The ADU-Ambiguity dataset is designed to evaluate the audio dialogue ambiguity handling ability from different phonetic elements, including intonation-based, pause-based, homophone-based, and repetition-based ambiguity. Notably, we firstly analyze the ambiguity within audio dialogues, specifically addressing the challenge of different intentions that share the same literal sentence, such as the word “Really!?” spoken with different intonations. In total, ADU-Bench comprises over 20,000 open-ended audio dialogues specifically designed for LALMs. An overall example of ADU-Bench is shown in Fig. 1.
@@
We benchmark 13 popular LALMs on our ADU-Bench and analyze the results. Our analysis reveals: (1) There is still considerable room for improvement in the audio dialogue understanding of existing open-sourced LALMs. (2) LALMs face challenges when dealing with skills, such as Mathematics and Coding, which involve mathematical symbols and formulas. (3) LALMs exhibit limitations in handling tasks related to Common Sense and Roleplay, as they lack a deeper understanding of human behavior. (4) Existing LALMs struggle to comprehend different meanings of audio dialogues that have the same transcriptions, but differ in phonetic elements, such as intonations, pause positions, and homophones. The evaluation code, datasets, and an open leaderboard will be made publicly available soon.
@@ 2 Related Work
Large Audio-language Models. Large audio-language models (LALMs) typically integrate audio modalities into large language models (LLMs) to extend their capabilities for general-purpose audio and language understanding. LALMs can be broadly classified into two types: end-to-end LALMs and cascaded LALMs. End-to-end LALMs can be further divided into two categories: (1) End-to-end LALMs specialize in audio understanding, which focus on integrating audio modality into LLMs, such as SpeechGPT (Zhang et al., 2023), BLSP (Wang et al., 2023), SALMONN (Tang et al., 2024), Qwen-Audio (Chu et al., 2023), and Mini-Omni (Xie and Wu, 2024). (2) End-to-end LALMs extend their capabilities beyond audio understanding, which align various modalities into a single LLM, such as PandaGPT (Su et al., 2023) and NExT-GPT (Wu et al., 2024). Another approach involves cascaded LALMs like the combination of an automatic speech recognition model, such as Whisper-large (Radford et al., 2023), and an LLM, such as GPT-4 (Achiam et al., 2023), to process a wide range of audio types. Our ADU-Bench aims to evaluate their performance in audio dialogue understanding across different domains.
Benchmarks for LALMs. As various LALMs have emerged, several benchmarks have been developed to evaluate their performance. Specifically, Dynamic-SUPERB (Huang et al., 2024) is the first dynamic and collaborative benchmark for evaluating instruction-tuning speech models and it primarily focuses on human speech processing. Another benchmark, AIR-Bench (Yang et al., 2024), is specifically designed for evaluating LALMs, encompassing multiple audio fundamental tasks and audio question answering. In the latter, LALMs are presented with an audio clip and a related question, requiring them to analyze the audio content and accurately answer the question. However, both benchmarks don’t assess the more practical open-ended audio dialogue understanding capabilities of LALMs. To bridge this gap, we propose ADU-Bench, which concentrates on evaluating LALMs in audio dialogue scenarios.
3 Data Collection and Statistics
@@
      ADU-Skill       Mathematics, Physics              GSM8K, MATH       WizardLM, ShareGPT       MBPP, MMLU       HotpotQA, StrategyQA       3,725
@@
      ADU-Multilingual       Arabic, Chinese, English              Alpaca, NQ-Bench       WebGLM
@@
      ADU-Ambiguity       Intonation-based              Phonetics and phonology       books
@@ 5.1 Experimental Settings
To benchmark the audio dialogue understanding of existing LALMs, we evaluate 13 foundational models with audio understanding capabilities. These models include PandaGPT-7B (Su et al., 2023), NExT-GPT-7B (Wu et al., 2024), Qwen-Audio-7B (Chu et al., 2023), Qwen-Audio-Chat-7B (Chu et al., 2023), Mini-Omni-0.5B (Xie and Wu, 2024), SpeechGPT-7B (Zhang et al., 2023), SALMONN-7B (Tang et al., 2024), SALMONN-13B (Tang et al., 2024), BLSP-7B (Wang et al., 2023), Whisper-large-v3 (Radford et al., 2023) with LLaMA-2-7B-Chat (Touvron et al., 2023), with LLaMA-3-8B-Instruct (MetaAI, 2024), with LLaMA-3-70B-Instruct (MetaAI, 2024), and with GPT-4-0613 (Achiam et al., 2023). Unless stated otherwise, the hyperparameters and setups used during the evaluation process remain consistent with those specified in the original papers of the respective models. For evaluation, we obtain two evaluation scores by swapping references and responses in the prompts for the GPT-4 evaluator and finally report the average scores for each model in Table 2. In addition, to avoid the bias of evaluation only using GPT-4, we apply various open-sourced LLMs for such evaluations, including LLaMA-3-70B-Instruct (MetaAI, 2024) and Qwen-2-72B-Instruct (Chu et al., 2023). More details about experimental settings are shown in Appendix C.
@@
Table 2: The average evaluation scores for audio dialogue understanding under 13 LALMs in our ADU-Bench.
Models Size ADU-Bench Average
General Skill Multilingual Ambiguity
PandaGPT 7B 1.02 0.98 0.98 0.50 0.87
NExT-GPT 7B 1.07 1.03 1.02 0.52 0.91
Qwen-Audio 7B 1.32 1.08 1.07 0.61 1.02
Mini-Omni 0.5B 2.31 1.96 1.55 1.67 1.87
SALMONN 7B 2.47 2.01 1.83 1.73 2.01
Qwen-Audio-Chat 7B 2.34 2.46 1.58 1.93 2.08
SpeechGPT 7B 3.99 3.56 1.42 2.25 2.81
SALMONN 13B 4.07 3.12 3.25 1.86 3.08
BLSP 7B 4.66 4.49 2.89 3.37 3.85
Whisper+LLaMA-2 7B 6.30 6.26 4.92 4.39 5.47
Whisper+LLaMA-3 8B 6.94 7.88 6.27 4.92 6.50
Whisper+LLaMA-3 70B 7.26 8.03 6.12 5.13 6.64
Whisper+GPT-4 - 8.42 8.62 8.07 5.54 7.66
5.2 Overall Results
@@
We report the experimental results for the performance of 13 different LALMs on audio dialogue understanding in Table 2 and provide a comprehensive analysis of them. Firstly, it can be observed that PandaGPT, NExT-GPT, and Qwen-Audio exhibit the lowest performances, with an average score value of about 1.00. It illustrates that although PandaGPT and NExT-GPT are end-to-end LALMs capable of processing a wide range of modalities, their performances on audio dialogue understanding are relatively lower. As for Qwen-Audio, a pre-trained base LALM, its weak capabilities in audio dialogue indicate a potential necessity for more specialized training to enhance its understanding of audio dialogues.
@@
Compared to them, Mini-Omni-0.5B, SALMONN-7B and Qwen-Audio-Chat show somewhat superior performance. This can be attributed to the fact that Mini-Omni-0.5B, SALMONN-7B, and Qwen-Audio-Chat have been developed under audio instruction tuning, making them suitable for a variety of audio-oriented scenarios. Moreover, SpeechGPT, SALMONN-13B, and BLSP have demonstrated even higher proficiency, as reflected in their average scores all about or exceeding 3.00. Among these, BLSP stands out with the highest average score of 3.85 among all LALMs. As SALMONN increases in size from 7B to 13B, its audio dialogue understanding capabilities also show improvement. In addition, both SpeechGPT and BLSP enable audio dialogue with LLMs using speech and exhibit impressive dialogue capabilities. Therefore, it can achieve enhanced performance when subjected to the targeted audio dialogue tuning for end-to-end LALMs, which highlights the importance of developing training strategies to improve audio dialogue understanding capabilities.
@@
Furthermore, cascaded LALMs, including LLaMA-2-7B, LLaMA-3-8B, LLaMA-3-70B, and GPT-4 with a Whisper model, obtain higher scores in audio dialogue understanding. Therein, GPT-4 leads the pack with the highest score of 7.66, which indicates that it is the best-performing model among the evaluated LALMs. Following it, LLaMA-3 (including LLaMA-3-8B and LLaMA-3-70B) ranks second, outperforming its predecessor, LLaMA-2. The improved performance of LLaMA-3 to LLaMA-2 highlights the effectiveness of updates in the LLaMA series.
In summary, cascaded LALMs outperform end-to-end LALMs with the specializing audio instruction tuning, which in turn surpass end-to-end LALMs with multi-modal understanding beyond audio, which indicates that further modality alignment should be developed for the open-ended audio dialogue understanding.
@@
Figure 2: The average scores across each domain for 4 datasets within ADU-Bench under 13 LALMs.

### v2

@@ Abstract
Large Audio-Language Models (LALMs), such as GPT-4o, have recently unlocked audio dialogue capabilities, enabling direct spoken exchanges with humans. The potential of LALMs broadens their applicability across a wide range of practical scenarios supported by audio dialogues. However, given these advancements, a comprehensive benchmark to evaluate the performance of LALMs in the open-ended audio dialogue understanding remains absent currently. To address this gap, we propose an Audio Dialogue Understanding Benchmark (ADU-Bench), which consists of 4 benchmark datasets. They assess the open-ended audio dialogue ability for LALMs in 3 general scenarios, 12 skills, 9 multilingual languages, and 4 categories of ambiguity handling. Notably, we firstly propose the evaluation of ambiguity handling in audio dialogues that expresses different intentions beyond the same literal meaning of sentences, e.g., “Really!?” with different intonations. In summary, ADU-Bench includes over 20,000 open-ended audio dialogues for the assessment of LALMs. Through extensive experiments on 16 LALMs, our analysis reveals that existing LALMs struggle with mathematical symbols and formulas, understanding human behavior such as roleplay, comprehending multiple languages, and handling audio dialogue ambiguities from different phonetic elements, such as intonations, pause positions, and homophones. The benchmark is available at https://adu-bench.github.io/.
@@ 1 Introduction
Large Audio-Language Models (LALMs) (Chu et al., 2023; Tang et al., 2024; Wu et al., 2023; Kong et al., 2024; Lin et al., 2024; Xie and Wu, 2024; Fu et al., 2024) have received attention for their abilities to handle various audio-related tasks. In particular, LALMs recently unlock unprecedented capabilities for interactive audio dialogues with humans. These dialogues are defined as a direct exchange of spoken language between LALMs and humans, which fosters a more dynamic mode of communication. Recent advances, such as GPT-4o (OpenAI, 2024), have enabled LALMs to engage in back-and-forth dialogues with humans and can observe various audio characteristics, which broadens their applicability across diverse real-world situations that rely on interactive audio dialogues.
@@
However, given these advancements, there is currently no comprehensive benchmark to evaluate LALMs’ performance in handling open-ended audio dialogue understanding. Previous benchmarks on LALMs predominantly focus on their performance in multiple fundamental tasks (Huang et al., 2024b, a), audio question answering with text-based instructions (Yang et al., 2024; Deshmukh et al., 2024; Sakshi et al., 2024; Wang et al., 2025) or audio dialogues in general scenarios (Ao et al., 2024; Chen et al., 2024). The absence of a comprehensive benchmark for evaluating LALMs in open-ended audio dialogues has led to suboptimal comparisons between different LALMs.
Open-ended audio dialogues, where users can directly engage with LALMs through audio, constitute a significant portion of real-world interactions. These dialogues can encompass many subjects, such as helpful and daily questions, domain-specific skills, and multiple different languages. Additionally, the variations in intonations or pause positions can allow speakers to express different intentions beyond the same literal meaning of sentences, adding further complexity to the dialogues. Therefore, the ability to handle open-ended audio dialogues effectively is crucial for LALMs to be truly useful in real-world applications.
In this work, we propose an Audio Dialogue Understanding Benchmark (ADU-Bench), a benchmark to evaluate the open-ended audio dialogue understanding for LALMs, which comprises 4 benchmark datasets as follows. (1) The ADU-General dataset assesses the general dialogue understanding of LALMs, including 3 scenarios, i.e., helpful questions to query search engines, daily questions happening among human dialogues, and daily statements without rich contexts. (2) The ADU-Skill dataset evaluates the skill-based dialogue ability, encompassing 12 different skills such as mathematics, physics, coding, etc. (3) The ADU-Multilingual dataset tests the multilingual dialogue understanding, covering 9 languages, including English, French, and Chinese, etc. (4) The ADU-Ambiguity dataset is designed to evaluate the audio dialogue ambiguity handling ability from different phonetic elements, including intonation-based, pause-based, homophone-based, and repetition-based ambiguity. Notably, we firstly analyze the ambiguity within audio dialogues, specifically addressing the challenge of different intentions that share the same literal sentence, such as the word “Really!?” spoken with different intonations. In total, ADU-Bench comprises over 20,000 open-ended audio dialogues for LALMs. An overall example of ADU-Bench is shown in Fig. 1.
@@
We benchmark 16 popular LALMs on our ADU-Bench and analyze the results. Our analysis reveals: (1) There is still considerable room for improvement in the audio dialogue understanding of existing open-sourced LALMs. (2) LALMs face challenges when dealing with skills, such as Mathematics and Coding, which involve mathematical symbols and formulas. (3) LALMs exhibit limitations in handling tasks related to Common Sense and Roleplay, as they lack a deeper understanding of human behavior. (4) Existing LALMs struggle to comprehend different meanings of audio dialogues that have the same transcriptions, but differ in phonetic elements, such as intonations, pause positions, and homophones. We include some demonstrations of our audio dialogues on our project page.
@@ 2 Related Work
Large Audio-language Models. Large audio-language models (LALMs) (Chu et al., 2023; Tang et al., 2024; Wu et al., 2023; Kong et al., 2024) typically integrate audio modalities into large language models (LLMs) (Touvron et al., 2023; OpenAI, 2024; Gao et al., 2024c, b, a; Fang et al., 2025; Zou et al., 2025; Kong et al., 2025) to extend their capabilities for general-purpose audio and language understanding. LALMs can be broadly classified into two types: end-to-end LALMs and cascaded LALMs. End-to-end LALMs can be further divided into two categories: (1) End-to-end LALMs specialize in audio understanding, which focus on integrating audio modality into LLMs, such as SpeechGPT (Zhang et al., 2023), BLSP (Wang et al., 2023), SALMONN (Tang et al., 2024), and Qwen-Audio (Chu et al., 2023). (2) End-to-end LALMs extend their capabilities beyond audio understanding, which align various modalities into a single LLM, such as PandaGPT (Su et al., 2023) and NExT-GPT (Wu et al., 2024). Another approach involves cascaded LALMs like the combination of an automatic speech recognition model, such as Whisper-large (Radford et al., 2023), and an LLM, such as GPT-4 (Achiam et al., 2023), to process a wide range of audio types. Our ADU-Bench aims to evaluate their performance in audio dialogue understanding across different domains.
Benchmarks for LALMs. Existing benchmarks for audio-related tasks can be broadly categorized into three areas: (1) fundamental audio tasks, (2) audio question answering with text-based instructions, and (3) audio dialogues. For benchmarks focusing on fundamental audio tasks (Huang et al., 2024b, a), evaluations are typically centered around specific objectives such as speech-to-text translation or emotion recognition. In audio question answering with text-based instructions (Yang et al., 2024; Deshmukh et al., 2024; Sakshi et al., 2024; Wang et al., 2025), models are required to interpret input audio and respond to input text-based instructions. In contrast, benchmarks for audio dialogues evaluate models to directly respond to audio queries without text-based instructions. While several established benchmarks (Ao et al., 2024; Chen et al., 2024) exist for audio dialogues, they predominantly focus on general scenarios, leaving a comprehensive benchmark unexplored. To bridge this gap, we propose ADU-Bench, which concentrates on evaluating LALMs in a wide range of audio dialogue scenarios.
3 ADU-Bench
@@
   ADU-Skill    Mathematics, Physics
   GSM8K, MATH
   WizardLM, ShareGPT
   MBPP, MMLU
   HotpotQA, StrategyQA
   3,725
@@
   ADU-Multilingual    Arabic, Chinese, English
   Alpaca, NQ-Bench
   WebGLM
@@
   ADU-Ambiguity    Intonation-based
   Phonetics and phonology
   books
@@ 5.1 Experimental Settings
To benchmark the audio dialogue understanding of existing LALMs, we evaluate 16 foundational models with audio understanding capabilities. These models include PandaGPT-7B (Su et al., 2023), NExT-GPT-7B (Wu et al., 2024), Qwen-Audio-7B (Chu et al., 2023), Qwen-Audio-Chat-7B (Chu et al., 2023), Mini-Omni-0.5B (Xie and Wu, 2024), SpeechGPT-7B (Zhang et al., 2023), Moshi-7B (Défossez et al., 2024), SALMONN-7B (Tang et al., 2024), SALMONN-13B (Tang et al., 2024), BLSP-7B (Wang et al., 2023), Step-Audio-Chat-130B (Huang et al., 2025), Whisper-large-v3 (Radford et al., 2023) with LLaMA-2-7B-Chat (Touvron et al., 2023), with LLaMA-3-8B-Instruct (MetaAI, 2024), with LLaMA-3-70B-Instruct (MetaAI, 2024), with GPT-4 (gpt-4o-0613) (Achiam et al., 2023), and GPT-4o (gpt-4o-audio-preview-2024-12-17) (OpenAI, 2024). Unless stated otherwise, the hyperparameters and setups used during the evaluation process remain consistent with those specified in the original papers of the respective models. For evaluation, we obtain two evaluation scores by swapping references and responses in the prompts for the GPT-4 evaluator and finally report the average scores for each model in Table 2. In addition, to avoid the bias of evaluation only using GPT-4, we apply various open-sourced LLMs for such evaluations, including LLaMA-3-70B-Instruct (MetaAI, 2024) and Qwen-2-72B-Instruct (Chu et al., 2023). In addition, we conduct a direct human evaluation on randomly selected 140 audio dialogues. Each sample is assessed by three human testers, who rate the generated responses. More details about experimental settings and human evaluation are in Appendix C and Appendix D, respectively.
@@
Table 2: The average evaluation scores for audio dialogue understanding under 16 LALMs in our ADU-Bench.
Models Size ADU-Bench Average Human
General Skill Multilingual Ambiguity Evaluation
PandaGPT 7B 1.02 0.98 0.98 0.50 0.87 -
NExT-GPT 7B 1.07 1.03 1.02 0.52 0.91 -
Qwen-Audio 7B 1.32 1.08 1.07 0.61 1.02 -
Mini-Omni 0.5B 2.31 1.96 1.55 1.67 1.87 -
SALMONN 7B 2.47 2.01 1.83 1.73 2.01 -
Qwen-Audio-Chat 7B 2.34 2.46 1.58 1.93 2.08 -
SpeechGPT 7B 3.99 3.56 1.42 2.25 2.81 -
Moshi 7B 4.37 3.08 1.49 2.81 2.94 -
SALMONN 13B 4.07 3.12 3.25 1.86 3.08 -
BLSP 7B 4.66 4.49 2.89 3.37 3.85 -
Step-Audio-Chat 130B 6.37 7.31 2.45 4.72 5.21 -
Whisper+LLaMA-2 7B 6.30 6.26 4.92 4.39 5.47 6.43
Whisper+LLaMA-3 8B 6.94 7.88 6.27 4.92 6.50 6.85
Whisper+LLaMA-3 70B 7.26 8.03 6.12 5.13 6.64 7.46
Whisper+GPT-4 - 8.42 8.62 8.07 5.54 7.66 8.02
GPT-4o - 8.64 8.97 8.16 6.87 8.16 8.58
5.2 Main Results
@@
We report the experimental results for the performance of 16 different LALMs on audio dialogue understanding in Table 2 and provide a comprehensive analysis of them. Firstly, it can be observed that PandaGPT, NExT-GPT, and Qwen-Audio exhibit the lowest performances, with an average score value of about 1.00. It illustrates that although PandaGPT and NExT-GPT are end-to-end LALMs capable of processing a wide range of modalities, their performances on audio dialogue understanding are relatively lower. As for Qwen-Audio, a pre-trained base LALM, its weak capabilities in audio dialogue indicate a potential necessity for more specialized training to enhance its understanding of audio dialogues.
@@
Compared to them, Mini-Omni-0.5B, SALMONN-7B and Qwen-Audio-Chat show somewhat superior performance. This can be attributed to the fact that Mini-Omni-0.5B, SALMONN-7B, and Qwen-Audio-Chat have been developed under audio instruction tuning, making them suitable for a variety of audio-oriented scenarios. Moreover, SpeechGPT, Moshi, SALMONN-13B, BLSP, and Step-Audio-Chat have demonstrated even higher proficiency, as reflected in their average scores all about or exceeding 3.00. Among these, BLSP stands out with the highest average score of 3.85 among all LALMs. As SALMONN increases in size from 7B to 13B, its audio dialogue understanding capabilities also show improvement. In addition, both SpeechGPT and BLSP enable audio dialogue with LLMs using speech and exhibit impressive dialogue capabilities. Therefore, it can achieve enhanced performance when using the targeted audio dialogue tuning for end-to-end LALMs.
@@
Furthermore, cascaded LALMs, including LLaMA-2-7B, LLaMA-3-8B, LLaMA-3-70B, and GPT-4 with a Whisper model, obtain higher scores in audio dialogue understanding. Therein, GPT-4 leads the pack with a high score of 7.66. Following it, LLaMA-3 (including LLaMA-3-8B and LLaMA-3-70B) ranks second, outperforming its predecessor, LLaMA-2. The improved performance of LLaMA-3 to LLaMA-2 highlights the effectiveness of updates in the LLaMA series.
Notably, for the advanced proprietary LALM, GPT-4o, achieves the highest average score of 8.16, which indicates that it is the best-performing model among the evaluated LALMs.
In addition, experimental results reveal that the GPT-4 evaluator demonstrates a significantly higher correlation with human evaluations, as shown in Table 2. Furthermore, we also conduct another human evaluation study, detailed in Section 5.4. These human evaluations verify the alignment between GPT-4 evaluator and human judgments.
@@
Figure 2: The average scores across each domain for 4 datasets within ADU-Bench under 16 LALMs.